OpenStudy (anonymous):

How do I cancel fractions when converting from point-slope form to standard form? For example, I'm currently tackling converting: y-0=-4/5 (x - -4) I end up with: y-0=-4/5x -16/5 but I'm confused on how to cancel the fractions? Thanks for your help!

OpenStudy (homeworksucks):

If you mean converting y=mx+b or y-1=m(x-1) into ax+by+c=0, you need to multiply by a common denominator